OpenStudy (anonymous):

How would I simpify fraction: 10/sqrt5?

zepdrix (zepdrix):

To simplify it, we want to multiply it by a fraction that will not change it's VALUE, but will change the way it looks. We can do this by multiplying by the number 1. But not just any 1, a fraction which is equivalent to 1.\[\large \frac{2}{2}=1\]Right? That's the idea we want to apply here. We'll multiply our fraction by,\[\large \frac{\sqrt5}{\sqrt5}\]

zepdrix (zepdrix):

\[\huge \frac{10}{\sqrt5}\cdot\frac{\sqrt5}{\sqrt5}\]What do you get when you multiply this out? :)

OpenStudy (anonymous):

22/5?

zepdrix (zepdrix):

The bottom of your fraction looks good. But the top unfortunately, we can't simplify any further. we have to leave the multiplication on top. But we do have one more step we can make. Here is what we have so far,\[\huge \frac{10\sqrt5}{5}\]

zepdrix (zepdrix):

The 10/5 we should be able to simplify a little bit.

OpenStudy (anonymous):

We simplify it to 2sqrt5/1?

zepdrix (zepdrix):

Yes very good!! :) And your final step would be to get rid of the 1 on the bottom. No need to write a fraction over 1.\[\huge \frac{2\sqrt5}{1}\quad = \quad 2\sqrt5\]Ok good job! :D

OpenStudy (anonymous):

Thanks so much!